production 13 gees of C2H2 requires consumption of how many g of H2O


1
Expert's answer
2020-10-10T08:36:11-0400

C2H2+2H2O=Ca(OH)2+C2H2

"n=\frac{m}{M}=\frac{13}{26}=0.5" mol

"\frac{x}{2}=\frac{0.5}{1}"

x=1

"m=n\times M=1*18=18" g